OpenStudy (anonymous):

The number of bacteria in a culture is modeled by the function n(t) = 600e^0.5t where t is measured in hours. (a) What is the initial number of bacteria? (b) What is the relative rate of growth of this bacterium population? Express your answer as a percentage. (c) How many bacteria are in the culture after 3 hours? (Round your answer to the nearest whole number.) (d) After how many hours will the number of bacteria reach 40,000? (Round your answer to two decimal places.)

OpenStudy (anonymous):

a) fot the initial number of bacteria u must think at hour zero n(t) = 600e^0.5(0)=600 b) for the rate of change u must think in a derivative dn/dt=300e^0.5t c) n(t) = 600e^0.5(3)=2689 d) Solve an equation 40000=600e^0.5t 66.67=e^0.5t ln(66.67)=0.5t t=8.4 == 8 hours b) im not sure because the change is not linear im not sure you cant express on percentage